# Ill-posedness of the Dirichlet problem for 2D Lagrangian mean curvature equation

### v1

@@ 3. Stage Proposition
Note
$\int_{\Omega}|\nabla\phi|^{2}\mathrm{d}x=-\int_{\Omega}\phi\Delta\phi\mathrm{d}x=-\int_{\Omega}\phi u_{1}\mathrm{d}x\leq\|\phi\|_{L^{2}}\|u_{1}\|_{L^{2}},$
@@
here we apply (3.32) in last inequality. By Theorem 8.15 in [12], we know
@@
Denote $D=\|u_{1}\|_{0}$ and $B=\|\zeta\|_{0}$. We consider the following two cases.
@@
and for any $\sigma\in(0,1)$, by the $C^{1,\sigma}$ estimates ((4.45)-(4.46) in [12]) ,
@@
It is remained to estimates $\mathcal{E}_{15}$. From (3.26), for $j=0,1$, we can derive
@@
Step 3.Adding the second rank-one matrix. We also mollify $(v^{\lozenge},w^{\lozenge})$ to get $(\tilde{v}^{\lozenge},\tilde{w}^{\lozenge})$ at length-scale $\lambda^{1-2\tau}$. Then by (3.28) and Lemma A.1, one also has
@@
dist(supp $\textit{dist}(\textrm{supp }a,\textrm{supp }\rho)=\lambda^{-\tau}.$
Furthermore, same as Step 2, using (3.16),(3.15) and (3.28), one is able to derive
@@
$\mathcal{E}_{2}:$ $=D(v^{\star},w^{\star})-(D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2})=\sum_{k=1}^{5}\mathcal{E}_{2k}$
@@
$\mathcal{E}_{21}:=$ $\lambda^{1-2\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\textrm{sym}(\nabla\Phi_{2}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}^{\lozenge}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
$\mathcal{E}_{22}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla\Phi_{1}),$
@@
$\mathcal{E}_{25}:=$ $((v^{\star}-v^{\lozenge})\cot\Theta)\cdot\mathrm{Id}.$
@@
In fact, when bounding the $C^{1}$ norm of the second matrix error, we only need to keep in mind that $\lambda^{2\tau-1}$(and $\lambda^{\tau}$) plays the same role as $\lambda^{\tau}$(and $\lambda$ repsectively) in Step 2. Therefore, the final matrix error can be calculated as follows
@@
As discussed in the beginning of Section 3, we show Theorem 1.1 (2) by seeking infinitely many symmetric $2\times 2$ matrixes $A$’s and $C^{1,\alpha}(\alpha<\frac{1}{5})$ functions $v$’s such that
@@
where $V$ solves the following Dirichlet problem:
(4.1) $\Delta V&=F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta,\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
We will solve the first system about $A$ together with a subsolution by seeking solutions to some Possion equations and the second system by constructing a sequence of adapted subsolutions via convex integration.
@@
Step 1. Matrix and initial subsolutions. We construct matrix $A$ and initial subsolutions at the same time. Let $u$ be the solution to the following Dirichlet problem,
$\Delta u&=0,\mbox{ in }\Omega,\\ u&=g,\mbox{ on }\partial\Omega,\end{aligned}\right.$
@@
$-\Delta\psi&=1-\det\nabla^{2}u,\mbox{ in }\Omega,\\ \psi&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
@@
$\Delta U&=F(u):=2\nabla u\cdot\nabla\cot\Theta+u\Delta\cot\Theta,\mbox{ in }\Omega,\\ U&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
@@
With $\Delta u=0$, we know $01-\det\nabla^{2}u=1+(\partial^{2}_{11}u)^{2}+(\partial^{2}_{12}u)^{2}>$ 0, and hence by the strong maximum principle, we have
@@
Note that $(v_{0},w_{0})$ and $A$ are not unique. For example, taking $v_{0}$ as before and $w_{0}\in\operatorname{\mathbb{R}}^{2}$ as arbitrary constant vector will also work.
Step 2. Iteration. With Proposition 3.1, for any $q\in\operatorname{\mathbb{N}}$, we construct inductively $(v_{q},w_{q},\rho_{q},H_{q})$ such that
@@
    If $x\in\overline{\Omega}\setminus\tilde{\Omega}_{q}$, then $(v_{q},w_{q},\rho_{q},H_{q})(x)=(v,w,\rho,H)(x).$
@@
    If $x\in\tilde{\Omega}_{j}$ for some $j\geq q,$ we get the following estimates
@@
Here the amplitude sequence $\delta_{q}$ and the frequency sequence $\lambda_{q}$ are inductively defined by
$\delta_{1}=\max_{x\in\overline{\Omega}}\rho^{2},\quad\lambda_{q+1}=\lambda_{q}^{b},\quad\lambda_{q}=M\delta_{q}^{-\frac{1}{2\beta}},$
where arbitrary constant $\beta\in(0,\frac{1}{5}),$
(4.3) $b=1+\frac{9\sigma}{1-5\beta}\quad(\text{then }5\beta(b-1)+9\sigma=b-1).$
and $M$ is a large constant to be fixed.
Indeed, $(v_{0},w_{0},\rho_{0},H_{0})$ has been constructed in Step 1 and it is not hard to verify $(1)_{0}-(4)_{0}$. For $q\geq 1,$ using Proposition 3.1 and following Step 2 in the proof of Proposition 2.1 of [4], we are able to inductively get $(v_{q},w_{q},\rho_{q},H_{q}).$ Since the proof is exactly same, we omit the details.
Step 3. Taking limit. By $(3)_{q}$ and $(5)_{q}$, we know $(\rho_{q},H_{q})$ is a Cauchy sequence in $C^{0}(\overline{\Omega})$, and $(v_{q},w_{q})$ is a Cauchy sequence in $C^{1}(\overline{\Omega})$, respectively. Hence we can assume
@@
Moreover, for any $\alpha\in(0,\frac{1}{5})$ and $\varepsilon>0$, we can choose $\beta\in(\alpha,\frac{1}{5})$ in Step 2, by $(4)_{q}-(5)_{q}$, $(4)_{q+1}$ and interpolation, we have for any $\alpha\in(0,\frac{\beta}{b^{2}})$,
@@
which implies $(v_{q},w_{q})$ uniformly converges to $(\bar{v},\bar{w})$ in $C^{1,\alpha}(\overline{\Omega})$. Since $\alpha$ and $b$ can be taken arbitrarily close to 00 and $1$ by (4.3), respectively, $(\bar{v},\bar{w})\in C^{1,\alpha}({\Omega}),$ for any $\alpha<\beta$. From $(5)_{q}$, for any $\epsilon>0,$ we also have
@@
Step 4. Very weak solution. We will show $\bar{v}$ is a very weak solution to (1.1). On the one hand, by construction, it is easy to see $(\bar{v},\bar{w})|_{\partial\Omega}=(v_{0},w_{0})|_{\partial\Omega}$. On the other hand, $\rho_{q}\to 0$ implies $A-D(\bar{v},\bar{w})=0.$ Thus for any $\phi\in C^{\infty}_{0}(\Omega)$, by integration by parts and curl crul $\mbox{curl crul }A=-1$, we have
$-\int_{\Omega}\phi\sin\Theta\mathrm{d}x=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot A)\mathrm{d}x$
$=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot\frac{1}{2}\nabla\bar{v}\otimes\nabla\bar{v})\mathrm{d}x-\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x,$
where we have used curl $\mbox{curl curl}(\mathrm{sym}\nabla\bar{w})=0$ and notation
@@
$0=\int_{\Omega}\partial_{1}\bar{v}\partial_{2}\bar{v}\partial_{12}^{2}(\phi\sin\Theta)$ $-\frac{1}{2}(\partial_{1}\bar{v})^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}\bar{v})^{2}\partial_{11}^{2}(\phi\sin\Theta)\mathrm{d}x$
$+\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x-\int_{\Omega}\phi\sin\Theta\mathrm{d}x.$
@@
In this appendix, we provide the sketch proofs of Lemmas 3.2-3.3 for completeness, which are as same as those in the first author [4]. We begin with the following two useful lemmas.
Lemma A.1 ([9]).
@@
with constant $C$ depending only on $s,r,\alpha,\phi$.
Lemma A.2 ([4]).
Let $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ be a simply connected open bounded set with smooth boundary and $H:\Sigma\to\operatorname{\mathbb{R}}^{2\times 2}$ a smooth $2\times 2$ symmetric positive definite matrix valued function, such that for some $10<\alpha<$ 1, and $c,M\geq 1$,
@@
Then there exists a smooth diffeomorphism $\Phi:\Sigma\to\operatorname{\mathbb{R}}^{2}$ and a smooth positive function $\vartheta:\Sigma\to\operatorname{\mathbb{R}}$ satisfying
@@
Sketch proof of Lemma 3.2.
Recall the definition of $\Gamma_{i}:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$, $i=1,2$,
@@
For any nonnegative integer $k$, it is not hard to show the following estimates.
@@
The derivatives of $\Gamma_{i}$, $i=1,2$ could be calculated directly by the chain rule, then the desired estimates are directly following by the assumptions on $\rho,H$, Lemma A.2 and (A.1). ∎
Sketch proof of Lemma 3.3.

### v2

@@ 3. The Stage Proposition
Noting
$\int_{\Omega}|\nabla\phi|^{2}dx=-\int_{\Omega}\phi\Delta\phi dx=-\int_{\Omega}\phi u_{1}dx\leq\|\phi\|_{L^{2}}\|u_{1}\|_{L^{2}},$
@@
Here we apply (3.32) in the last inequality. By Theorem 8.15 in [GT], we know
@@
Let $D=\|u_{1}\|_{0}$ and $B=\|\zeta\|_{0}$. We consider the following two cases.
@@
and for any $\sigma\in(0,1)$, by the $C^{1,\sigma}$ estimates ((4.45)-(4.46) in [GT]) ,
@@
It remains to estimates $\mathcal{E}_{15}$. From (3.26), for $j=0,1$, we can derive
@@
[…]
@@
$\textit{dist}(\textrm{supp }a,\textrm{supp }\rho)=\lambda^{-\tau}.$
Furthermore, similar to Step 2, using (3.16),(3.15) and (3.28), one is able to derive
@@
$\mathcal{E}_{2}:$ $=D(v^{\star},w^{\star})-\left[D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2}\right]=\sum_{k=1}^{5}\mathcal{E}_{2k}$
@@
$\mathcal{E}_{21}:=$ $\lambda^{1-2\tau}\left[\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\textrm{sym}(\nabla\Phi_{2}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}^{\lozenge}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)\right],$
$\mathcal{E}_{22}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}\left[(\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla a\right]+\partial_{t}\Gamma_{1}\textrm{sym}\left[(\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla\Phi_{1}\right],$
@@
$\mathcal{E}_{25}:=$ $\left[(v^{\star}-v^{\lozenge})\cot\Theta\right]\cdot\mathrm{Id}.$
@@
In fact, when bounding the $C^{1}$ norm of the second matrix error, we only need to keep in mind that $\lambda^{2\tau-1}$(and $\lambda^{\tau}$) plays the same role as $\lambda^{\tau}$(and $\lambda$ repsectively) in Step 2. Therefore, the final matrix error can be calculated as follows:
@@
As discussed at the beginning of Section 3, we show Theorem 1.1 (2) by seeking infinitely many symmetric $2\times 2$ matrixes $A$’s and $C^{1,\alpha}(\alpha<\frac{1}{5})$ functions $v$’s such that
@@
where $V$ satisfies the following Dirichlet problem:
(4.1) $\left\{\begin{aligned} \Delta V&=F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta,\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right$.
We will solve the first system about $A$ together with a subsolution by seeking solutions to some Poisson equations and the second system by constructing a sequence of adapted subsolutions via convex integration.
@@
Step 1. Matrix and initial subsolutions. We construct the matrix $A$ and initial subsolutions at the same time. Let $u$ be the solution to the following Dirichlet problem,
$\left\{\begin{aligned} \Delta u&=0,\mbox{ in }\Omega,\\ u&=g,\mbox{ on }\partial\Omega,\end{aligned}\right$.
@@
$\left\{\begin{aligned} -\Delta\psi&=1-\det\nabla^{2}u,\mbox{ in }\Omega,\\ \psi&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
@@
$\left\{\begin{aligned} \Delta U&=F(u):=2\nabla u\cdot\nabla\cot\Theta+u\Delta\cot\Theta,\mbox{ in }\Omega,\\ U&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
@@
With $\Delta u=0$, we know $1-\det\nabla^{2}u=1+(\partial^{2}_{11}u)^{2}+(\partial^{2}_{12}u)^{2}>0$, and hence by the strong maximum principle, we have
@@
Note that $(v_{0},w_{0})$ and $A$ are not unique. For example, taking $v_{0}$ as before and $w_{0}\in\operatorname{\mathbb{R}}^{2}$ as an arbitrary constant vector will also work.
Step 2. Iteration. We define the amplitude sequence $\delta_{q}$ and the frequency sequence $\lambda_{q}$ inductively by
$\delta_{1}=\max_{x\in\overline{\Omega}}\rho_{0}^{2},\quad\lambda_{q+1}=\lambda_{q}^{b},\quad\lambda_{q}=M\delta_{q}^{-\frac{1}{2\beta}},$
where $\beta$ is a arbitrary fixed constant in $(0,\frac{1}{5})$, and
(4.2) $b=1+\frac{9\sigma}{1-5\beta}\quad(\text{then }5\beta(b-1)+9\sigma=b-1),$
and $M$ is a large constant to be fixed. We then decomposes $\overline{\Omega}$ according to the level sets of $\rho_{0}$,
$\Omega_{q}=\left\{x\in\overline{\Omega}:\,\rho_{0}(x)\geq 2\delta_{q+2}^{1/2}\right\},\,\,\tilde{\Omega}_{q}=\left\{x\in\overline{\Omega}:\,\rho_{0}(x)\geq\frac{3}{2}\delta_{q+2}^{1/2}\right\}$
for any $q\in\mathbb{N}$. With Proposition 3.1, for any $q\in\operatorname{\mathbb{N}}$, we inductively construct $(v_{q},w_{q},\rho_{q},H_{q})$ such that the following hold:
@@
    If $x\in\overline{\Omega}\setminus\tilde{\Omega}_{q}$, then $(v_{q},w_{q},\rho_{q},H_{q})(x)=(v_{0},w_{0},\rho_{0},H_{0})(x).$
@@
    If $x\in\tilde{\Omega}_{j}$ for some $j\geq q,$ we get the following estimates:
@@
Indeed, $(v_{0},w_{0},\rho_{0},H_{0})$ has been constructed in Step 1 and it is not hard to verify $(1)_{0}-(4)_{0}$. Similar to Step 2 of the proof of Proposition 2.1 in [Cao], by using Proposition 3.1 we are able to inductively get $(v_{q},w_{q},\rho_{q},H_{q})$ for $q\geq 1$. Since the proof is exactly the same as in [Cao], we omit the details.
Step 3. Taking limits. By $(3)_{q}$ and $(5)_{q}$, we know $(\rho_{q},H_{q})$ is a Cauchy sequence in $C^{0}(\overline{\Omega})$, and $(v_{q},w_{q})$ is a Cauchy sequence in $C^{1}(\overline{\Omega})$, respectively. Hence we can assume
@@
Moreover, for any $\alpha\in(0,\frac{1}{5})$ and $\varepsilon>0$, we can choose $\beta\in(\alpha,\frac{1}{5})$ in Step 2. By $(4)_{q}-(5)_{q}$, $(4)_{q+1}$ and interpolation, we have for any $\alpha\in(0,\frac{\beta}{b^{2}})$,
@@
which implies $(v_{q},w_{q})$ uniformly converges to $(\bar{v},\bar{w})$ in $C^{1,\alpha}(\overline{\Omega})$. Since $\alpha$ and $b$ can be taken arbitrarily close to 0 and $1$ by (4.2), respectively, $(\bar{v},\bar{w})\in C^{1,\alpha}({\Omega}),$ for any $\alpha<\beta$. From $(5)_{q}$, for any $\epsilon>0,$ we also have
@@
Step 4. Very weak solutions. We will show $\bar{v}$ is a very weak solution to (1.1). On the one hand, by construction, it is easy to see $(\bar{v},\bar{w})|_{\partial\Omega}=(v_{0},w_{0})|_{\partial\Omega}$. On the other hand, $\rho_{q}\to 0$ implies $A-D(\bar{v},\bar{w})=0.$ Thus for any $\phi\in C^{\infty}_{0}(\Omega)$, by integration by parts and $\mbox{curl curl }A=-1$, we have
$-\int_{\Omega}\phi\sin\Theta dx=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot A)dx$
$=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot\frac{1}{2}\nabla\bar{v}\otimes\nabla\bar{v})dx-\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)dx,$
where we have used $\mbox{curl curl}(\mathrm{sym}\nabla\bar{w})=0$ and the notation
@@
$0=\int_{\Omega}\partial_{1}\bar{v}\partial_{2}\bar{v}\partial_{12}^{2}(\phi\sin\Theta)$ $-\frac{1}{2}(\partial_{1}\bar{v})^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}\bar{v})^{2}\partial_{11}^{2}(\phi\sin\Theta)dx$
$+\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)dx-\int_{\Omega}\phi\sin\Theta dx.$
@@
In this appendix, we provide the sketch proofs of Lemmas 3.2-3.3 for completeness, which are as same as those in [Cao]. We begin with the following two useful lemmas.
Lemma A.1 ([CDS12]).
@@
with the constant $C$ depending only on $s,r,\alpha,\phi$.
Lemma A.2 ([Cao]).
Let $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ be a simply connected open bounded set with a smooth boundary and $H:\Sigma\to\operatorname{\mathbb{R}}^{2\times 2}$ a smooth $2\times 2$ symmetric positive definite matrix-valued function, such that for some $0<\alpha<1$, and $c,M\geq 1$,
@@
Then there exist a smooth diffeomorphism $\Phi:\Sigma\to\operatorname{\mathbb{R}}^{2}$ and a smooth positive function $\vartheta:\Sigma\to\operatorname{\mathbb{R}}$ satisfying
@@
A sketch proof of Lemma 3.2.
Recall the definition of $\Gamma_{i}:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$, $i=1,2$:
@@
For any nonnegative integer $k$, it is not hard to show the following estimates:
@@
The derivatives of $\Gamma_{i}$, $i=1,2$ could be calculated directly by the chain rule, then the desired estimates directly follow from the assumptions on $\rho,H$, Lemma A.2 and (A.1). ∎
A sketch proof of Lemma 3.3.